• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Lindgren

    Log entry time 16:05:06 on June 8,2010

    Entry number 325298

    keyword=summary

    Day Shift Starting 8:00 AM No Beam
    SL R. Lindgren
    TO Nicholas Lumbinskyu
    3rd Werner Boeglin

    Waiting for beam, MCC reports beam in afew minutes
    8:05 Beam back
    Redxuce to 5 ua Do a spot check. ok logged it in
    Beam ramping up to 50 ua
    Compton started at 7:40 AM
    8:08 Start Run 4403 50 ua beam back to production running
    8:30 Dithering was off stop run called jonathan to come in to start it
    Start Run 4404 Take data anyway until Jonathan comes in
    8;50 jonathan stated dithering following standard procedure no change in previous amplitudes of coils. Operation manuals is out fo date.
    keep running no change in run numbers
    9:03 Stop Run 4404

    Beam studies
    Going to control access
    Expect to start up in 2 hours

    Control Access
    Request to power down BigBox.
    Comunication error on BigBox controls
    Tried to reboot IOC using hareboot7
    BigBox remote controls won't reboot
    Rquired local control to bring magnet down
    Requires radiation watch
    Zafa Ahmad volunteers to watch at 9:40 AM

    10;00 to 12:00
    Jack repairs BigBox controls
    Heidi repairs flow meter
    Albert checks cables for Bogdan


    13:00
    Beam startup procedure done
    Waiting for MCC to ramp up beam to 50 ua

    summary of day shift production runs

    run start stop charge

    4403 8:08 8:31 38.4 mC
    4404 8:40 9:03 39.2 mC
    4405 13:08 13:53 128.9 mC
    4406 13:53 14:44 125.3 mC
    4407 14:45 15:35 103.3 mC
    4408 15:35

    Day shift ends


    FIGURE 1